ln e^(log(x+1)+log(x-2)) = (1/2)e^(ln(2)), what is x?
Found 2 solutions by psbhowmick, lwsshak3:
Answer by psbhowmick(878)

Can you type the equation properly... it is not very clear which are log to the base 10 and which are log to the base e.

Answer by lwsshak3(11628)

ln e^(log(x+1)+log(x-2)) = (1/2)e^(ln(2)), what is x?
***
log of base raised to a power=power
ln e^(log(x+1)+log(x-2))=log(x+1)+log(x-2)=log(x+1)(x-2)
..
Base raised to the log of number=number
(1/2)e^(ln(2))=(1/2)*2=1
..
log(x+1)(x-2)=1
10^1=(x+1)(x-2)
10=x^2-x-2
x^2-x-12=0
(x-4)(x+3)=0
x=4
or
x=-3